# Quadratic Volatility from the Pöschl-Teller Potential and Hyperbolic Geometry

### v1

@@ I. Introduction
Operational protocols for contingent claim valuation have long centered on the framework established by Black and Scholes [8] and Merton [34]. This paradigm, predicated on the principle of dynamic replication in a frictionless market, yields a unique pricing partial differential equation (PDE). A critical assumption embedded within this formulation is that the volatility of the underlying asset’s returns remains constant throughout the option’s tenor, implying that asset prices follow a geometric Brownian motion [8, 34]. The elegance of this construct is its analytical tractability; a sequence of canonical transformations recasts the pricing PDE into the heat equation, a fundamental equation of mathematical physics [5]. This equivalence hints at a deeper structure, suggesting that the financial pricing operator can be viewed as a Hamiltonian, and the valuation problem itself as a quantum mechanical system evolving in imaginary time [5].
This theoretical consistency, however, is confronted by a persistent empirical anomaly. Market-observed prices for European options, when inverted through the Black-Scholes formula, do not yield a constant volatility. Instead, the implied volatility parameter exhibits a distinct, systematic dependence on both strike price and time to maturity [21]. This phenomenon, widely documented as the ”volatility smile” or ”skew,” became particularly pronounced in equity index markets following the 1987 crash and represents a significant departure from the log-normal return distribution assumed by the standard model [40, 6]. The structure of the smile indicates that market participants assign a higher probability to large price movements—so-called ”fat tails”—than the Black-Scholes framework permits [11].
The existence of the volatility smile necessitates the development of models that relax the constant volatility assumption. One major line of inquiry led to the formulation of local volatility models, wherein volatility is a deterministic function of the asset price and time. The seminal works of Dupire [17] and Derman and Kani [14] demonstrated that it is possible to construct a unique risk-neutral diffusion process consistent with the complete surface of observed option prices. Alternative approaches have incorporated stochastic volatility and jump-diffusion processes to capture the smile and skew dynamics [7, 11]. While these models offer greater realism, they often sacrifice the analytical tractability of the original Black-Scholes equation.
@@ II. The Quadratic Normal Volatility Model
partitions the parameter space into three distinct regimes, each corresponding to a space of constant curvature [39, 48]. When $\Delta>0$, the polynomial has two distinct real roots, and the underlying manifold possesses a hyperbolic geometry, corresponding to a space of constant negative curvature. When $\Delta$=0, the polynomial has a single real root of multiplicity two, which corresponds to a Euclidean geometry, a space of zero curvature. Finally, when $\Delta$¡0, the polynomial has two complex conjugate roots, and the manifold exhibits a spherical geometry, corresponding to a space of constant positive curvature. This classification reveals that the choice of financial parameters (a,b,c) is not merely a curve-fitting exercise but is tantamount to selecting the fundamental geometry of the state space. The analytical solvability of the QNV model, which will be demonstrated in subsequent sections, is a direct consequence of the maximal symmetry inherent in these constant curvature spaces. The valuation PDE articulated in Equation (1) represents a general parabolic equation with state-dependent coefficients. Direct analytical resolution is obstructed by the non-constant diffusion term $\sigma(S)^{2}$ and the presence of a first-order spatial derivative. To render the problem amenable to spectral analysis, a sequence of canonical transformations is executed, a procedure designed to reduce the financial operator to a self-adjoint Sturm-Liouville form [31, 30]. This methodology is systematic and serves to diagonalize the pricing operator, thereby exposing its underlying spectral structure [20].
@@ Appendix A Exposition of the Canonical Transformation
$\frac{\partial C}{\partial t}+\frac{1}{2}\sigma(S)^{2}\frac{\partial^{2}C}{\partial S^{2}}+rS\frac{\partial C}{\partial S}-rC$ $=0.$
@@
$-e^{-r\tau}\left(\frac{\partial v}{\partial\tau}-rv\right)+\frac{1}{2}\sigma(S)^{2}e^{-r\tau}\frac{\partial^{2}v}{\partial S^{2}}+rSe^{-r\tau}\frac{\partial v}{\partial S}-re^{-r\tau}v$ $=0.$
@@ Appendix C Special Functions
$(1-z^{2})\frac{d^{2}w}{dz^{2}}-2z\frac{dw}{dz}+\left[\nu(\nu+1)-\frac{\mu^{2}}{1-z^{2}}\right]w$ $=0.$

### v2

@@ I. Introduction
Operational protocols for contingent claim valuation have long centered on the framework established by Black and Scholes [8] and Merton [34]. This paradigm, predicated on the principle of dynamic replication in a frictionless market, yields a unique pricing partial differential equation (PDE). A critical assumption embedded within this formulation is that the volatility of the underlying asset’s returns remains constant throughout the option’s tenor, implying that asset prices follow a geometric Brownian motion [8, 34]. The elegance of this construct is its analytical tractability; a sequence of canonical transformations recasts the pricing PDE into the heat equation, a fundamental equation of mathematical physics [5]. This equivalence hints at a deeper structure, suggesting that the financial pricing operator can be viewed as a Hamiltonian, and the valuation problem itself as a quantum mechanical system evolving in imaginary time [5]. This theoretical consistency, however, is confronted by a persistent empirical anomaly. Market-observed prices for European options, when inverted through the Black-Scholes formula, do not yield a constant volatility.
Instead, the implied volatility parameter exhibits a distinct, systematic dependence on both strike price and time to maturity [21]. This phenomenon, widely documented as the "volatility smile" or "skew," became particularly pronounced in equity index markets following the 1987 crash and represents a significant departure from the log-normal return distribution assumed by the standard model [40, 6]. The structure of the smile indicates that market participants assign a higher probability to large price movements—so-called "fat tails"—than the Black-Scholes framework permits [11]. The existence of the volatility smile necessitates the development of models that relax the constant volatility assumption. One major line of inquiry led to the formulation of local volatility models, wherein volatility is a deterministic function of the asset price and time. The seminal works of Dupire [17] and Derman and Kani [14] demonstrated that it is possible to construct a unique risk-neutral diffusion process consistent with the complete surface of observed option prices. Alternative approaches have incorporated stochastic volatility and jump-diffusion processes to capture the smile and skew dynamics [7, 11]. While these models offer greater realism, they often sacrifice the analytical tractability of the original Black-Scholes equation.
@@ II. The Quadratic Normal Volatility Model
partitions the parameter space into three distinct regimes, each corresponding to a space of constant curvature [39, 48]. When $\Delta>0$, the polynomial has two distinct real roots, and the underlying manifold possesses a hyperbolic geometry, corresponding to a space of constant negative curvature. When $\Delta$=0, the polynomial has a single real root of multiplicity two, which corresponds to a Euclidean geometry, a space of zero curvature. Finally, when $\Delta$<0, the polynomial has two complex conjugate roots, and the manifold exhibits a spherical geometry, corresponding to a space of constant positive curvature. This classification reveals that the choice of financial parameters (a,b,c) is not merely a curve-fitting exercise but is tantamount to selecting the fundamental geometry of the state space. The analytical solvability of the QNV model, which will be demonstrated in subsequent sections, is a direct consequence of the maximal symmetry inherent in these constant curvature spaces. The valuation PDE articulated in Equation (1) represents a general parabolic equation with state-dependent coefficients. Direct analytical resolution is obstructed by the non-constant diffusion term $\sigma(S)^{2}$ and the presence of a first-order spatial derivative. To render the problem amenable to spectral analysis, a sequence of canonical transformations is executed, a procedure designed to reduce the financial operator to a self-adjoint Sturm-Liouville form [31, 30]. This methodology is systematic and serves to diagonalize the pricing operator, thereby exposing its underlying spectral structure [20].
@@ Appendix A Exposition of the Canonical Transformation
$\frac{\partial C}{\partial t}+\frac{1}{2}\sigma(S)^{2}\frac{\partial^{2}C}{\partial S^{2}}+rS\frac{\partial C}{\partial S}-rC$ =0=0.= 0 .
@@
$-e^{-r\tau}\left(\frac{\partial v}{\partial\tau}-rv\right)+\frac{1}{2}\sigma(S)^{2}e^{-r\tau}\frac{\partial^{2}v}{\partial S^{2}}+rSe^{-r\tau}\frac{\partial v}{\partial S}-re^{-r\tau}v$ =0=0.= 0 .
@@ Appendix C Special Functions
$(1-z^{2})\frac{d^{2}w}{dz^{2}}-2z\frac{dw}{dz}+\left[\nu(\nu+1)-\frac{\mu^{2}}{1-z^{2}}\right]w$ =0=0.= 0 .